OpenStudy (anonymous):

Ill fan and medal ! find exact value of cos(a+b) under following conditions sin(a)=40/41 and sin(b)=15/17

OpenStudy (anonymous):

\[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\] you have \[\sin(a)=\frac{40}{41},\sin(b)=\frac{15}{17}\] and need to find \[\cos(a), \cos(b)\]